You are here: Foswiki>LabstaffWeb Web>LabstaffDowntime>20100502Cooling (03 May 2010, lindss2@LAB.CS.RPI.EDU)EditAttach
CII MRI cluster shut down due to cooling problems 2010-05-02.

Message from HVAC:
Saturday morning, one chiller went in alarm and tripped due to extended
surge, which means chiller was not loading up. Have to reset couple of
times to get it going.

Sunday, chillers tripped due to "High head pressure", (a) caused by  the
failure of (a) East sump make-up water valve, low water level in cooling
water sump and (b) tripped cooling tower fan for the west sump. There
were combination of faults, got resolved around 8:30PM or so.

Both chillers are now running satisfactory.

Diaphragm parts are on order for the makeup water valve and should be
here tomorrow morning. Meantime make up for the condenser water sump
will be handled by the west sump and keeping watch.

Service restored 20100503.
Topic revision: r2 - 03 May 2010, lindss2@LAB.CS.RPI.EDU
 

This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback